# Generalizable Human Gaussians from Single-View Image

### v4

@@ 3.4 Back-view Refinement
Back-view hallucination poses a significant challenge in single-view human reconstruction. As shown in Fig. 4, directly using diffusion models to generate a back-view image can result in incorrect perspective projection with the front-view image as well as unrealistic texture by sith24 . The reason is that their diffusion is conditioned on the back-view mask, and thus can only be applied for orthogonal projection where back-view mask can be directly flipped with the front-view image. However, the back-view mask is not available during our inference stage since our prediction is based on perspective projection. To address this issue, we adopt a generate-then-refine strategy that leverages the diffusion prior to produce a perspective-fitted and realistic back-view image that is suitable for the subsequent two-view fusion stage. We train a ControlNet control23 to enhances realistic details based on our coarse results as shown in Fig. 3 right part. Specifically, we generate coarse back-view rendering by our HGM for the training dataset and only train the ControlNet and keep the base Stable Diffusion model as fixed.
$\mathcal{L}_{CN}=\mathbb{E}_{\mathbf{z}_{0},t,\mathbf{c}_{t},\boldsymbol{\epsilon}_{t}\sim\mathcal{N}(0,1)}\left[\left\|\boldsymbol{\epsilon}-\epsilon_{\theta}(\mathbf{z}_{t},t,\mathbf{c}_{t},\mathbf{y})\right\|_{2}^{2}\right]$ (4)
where $y$ is the text prompt. We set it as ‘Best quality’ and the negtive prompt as ‘blur, bad anatomy, bad hands, cropped, worst quality’ during inference, $\mathbf{c}_{t}$ is the coarse back-view image from our HGM rendering $I_{B}^{c}$ , which is the ControlNet condition. We carefully design the reversing process by adding small amout of noise to the VAE encoded latent of $I_{B}^{c}$ to keep the original content as much as possible. The sampling process takes around 2 seconds. In Fig. 4, we showcase our generated and refined back-view results, comparing them with the back-view hallucination diffusion network in SiTH sifu24 . Our results maintain high resolution and generate details such as the hair for the first woman and the wrinkles in the clothes, whereas SiTH sifu24 produces artifacts and unrealistic hallucination results and are also not perspectively fitted to the input image.
@@ 3.5 Two-view Reconstruction
Specifically, our fusion HGM model retains the design of the coarse HGM model with the additional refined back-view image as the input. The coarse HGM and fusion HGM models are trained separately with ground-truth one view and two views as input, as well as ground-truth SMPL-X. The objective function for HGM training includes $\mathcal{L}_{2}$ color loss, $\mathcal{L}_{rgb}$, VGG-based LPIPS perceptual loss, $\mathcal{L}_{lpips}$ perceptual , and $\mathcal{L}_{2}$ background mask loss $\mathcal{L}_{bg}$ with ground truth masks. Each of these losses has corresponding weights that are treated as hyperparameters:
@@ 3.6 Implementation
Our model is trained on 4 NVIDIA RTX A6000 with batch size of 4 for 20 hours. Our input image size is 512×512, and the number of Gaussians for each view is 256×256, totaling 65,536 Gaussians per view. For SMPL-X estimation, we use PIXIEpixie21 . Network structures and more implementation details are in the appendix.
@@ 4 Experiments
We conduct experiments on the publicly available 3D human datasets THuman2.0 thuman21 , CustomHumanscustom23 and HuMMan ntuhuman22 . Our method is compared with state-of-the-art (SOTA) methods in both novel view synthesis and 3D mesh reconstruction. We train our HGM on 500 human scans from THuman2.0 dataset following sifu24 . We render the images with resolution of 512×512 and using weak perspective camera on 12 fixed cameras evenly distributed with the azimuths from 0 to 360 degree. During evaluation, all the methods are tested without the ground truth SMPL-X. We follow the train and test list from SIFU sifu24 and SHERF sherf23 to evaluate our method on THuman2.0 and HuMMan dataset. For CustomHumans dataset we use 45 scans for cross-dataset evaluation containing loosing clothes and challenging poses. For novel view synthesis, We use PSNR, SSIM, LPIPS as evaluation metrics. For 3D reconstruction, we use commonly used Chamfer Distance (CD), Point-to-Surface Distance (P2S), and Normal Consistency as the evaluation metrics.
@@ 4.1 Novel View Synthesis
SiTH sith24 17.12 0.843 0.155 18.09 0.856 0.144
LGM lgm24 18.34 0.856 0.134 19.87 0.877 0.132
SV3D sv3d24 19.11 0.892 0.117 20.86 0.902 0.112
SIFU sifu24 22.10 0.924 0.0794 20.83 0.898 0.117
@@
For novel view synthesis, we compare our method with mesh SOTA human reconstruction methods ECON econ23 , SIFU sifu24 and SiTH sith24 , as well as multiview diffusion reconstruction method LGM(finetuned with the same training data) lgm24 and video diffusion method SV3D sv3d24 on THuman2.0 thuman21 and CustomHuman custom23 . We also compare our method with state-of-the-art HumanNeRF methods: SHERF sherf23 , MPS-NeRF mps22 , and NHP nhp21 on the HuMMan dataset ntuhuman22 .
As shown in the Tab. 1 and Tab. 2, our method significantly surpasses state-of-the-art single-view human reconstruction methods in all evaluation metrics for the three datasets. As shown in Fig. 6, LGM lgm24 generates incorrect blue color and inconsistent content. SiTH sith24 fails to model loose clothes due to the high dependency of the SMPL-X model. Side views are blurry and unrealistic in SIFU’s results. SV3D sv3d24 generates strange colors and wrong human pose. Compared with these methods, ours generates more realistic and consistent rendering especially for the unseen regions such as clothes wrinkles and hair that are well-fitted to the front views and more robust to initial SMPL-X estimation errors thanks to our iterative refinement. We provide rendering videos in 360 degree comparison with other methods in the Appendix. We also provide a visual comparison with the SOTA NeRF-based method SHERF sherf23 on the HuMMan dataset in the appendix.
@@
NHP nhp21 18.99 0.845 0.182
MPS-NERF mps22 17.44 0.824 0.193
SHERF sherf23 20.83 0.891 0.125
@@
For mesh reconstruction we extract the 3D mesh by densely rendering the depth map with Gaussian render and using TSDFusion to extract the surface followed by a fast optimizaiton based on the normal map obtained in section3.3.2. We compare our results with SOTA human surface reconstruction methods ECON econ23 , SIFU sifu24 and SiTH sith24 . Note that our method do not use the 3D ground truth for supervision, but can also achieve best performance compared with all those fully supervised methods. Thanks to our interative SMPL-X refinement. Our method alleviates commonly occurring problems of bent legs and incorrect postures found in previous methods, as shown in Fig.5. Our HGM model reconstructs more accurate geometry with the prior learned from our dual branch Gaussian reconstruction model as well as our innovative SMPL-X refinement. We provide a qualitative 3D reconstruction comparison in Fig.7. As shown in the figure, ECON and SIFU suffer from bending legs and wrong arms problems. SiTH generates an over-smoothed surface and missing parts. Our method can reconstruct more accurate poses while preserving geometric details.
Table 3: 3D reconstruction comparison with SOTA methods.
Only our method trained without 3D supervision. THuman2.0 CustomHumans Method Chamfer $\downarrow$ P2S $\downarrow$ Normal $\uparrow$ Chamfer $\downarrow$ P2S $\downarrow$ NC $\uparrow$ ECON econ23 2.342 2.431 0.765 2.107 2.355 0.771 SiTH sith24 2.519 2.442 0.786 2.223 2.584 0.785 SIFU sifu24 2.063 2.205 0.792 1.864 1.976 0.778 Ours 2.134 2.118 0.823 1.729 1.835 0.834
@@ 4.3 Ablation Studies
We conduct ablation studies to evaluate the effectiveness of our SMPL-X dual branch Gaussian prediction model, coarse-to-fine refinement strategy, and back-view refine ControlNet, as well as our SMPL-X refinement. We show the quantitative ablation results in Table 4. The performance decrease when any component is removed. SMPL-X dual branch plays an important role in adding human priors through structural features to Image Gaussians predicted by UNet. We visualize the rendered images using our model without SMPL-X dual branch as the guidance and those produced by our full model, as shown in Fig. 8. Without SMPL-X dual branch as guidance, the side view images exhibit significant artifacts, such as misaligned arms and unnatural shapes of clothes and heads, highlighted in the red boxes. This demonstrates the effectiveness of our SMPL-X dual branch Gaussian prediction design.The predicted Gaussians lack human shape and pose priors without SMPL-X guidance, resulting in unnatural shapes and poses. The initial SMPL-X prediciton is not accurate, we ablates the effectiveness of our iterative SMPL-X refinement with the 3D prior learned in our HGM model. We also visualize the SMPL-X refinement in the appendix. Two-view refinement double the number of the Gaussians to improve the reconstruction quality. Gaussians tend to concentrate more on the front view without the two-view refinement strategy, leading to poorer rendering of the back part. Additionally, the diffusion-based refinement is crucial for improving the novel view synthesis quality, especially for the back-view images as shown in Fig. 4. The best performance is achieved with all three components.
4.4 Discussions
Our method improves upon previous approaches like ECON econ23 and SIFU sifu24 by leveraging our pre-trained HGM model to incorporate 3D priors, specifically side view masks, for enhanced SMPL-X refinement. Unlike earlier techniques that only align 2D front-view masks and normals, our approach achieves better reconstruction accuracy and alignment. The refined SMPL-X also benefits our human Gaussians reconstruction by providing structural information encoded in learnable tokens. Our 3D Gaussians-based method offers significant advantages in rendering speed, achieving 300 FPS compared to NeRF-based methods like SHERF sherf23 , which manages only 2 FPS. Furthermore, the mesh extracted from our 3D Gaussians with normal refinement attains high 3D reconstruction quality. In summary, our method excels in both high-quality rendering and accurate 3D reconstruction, offering a comprehensive solution.
@@ Appendix A Appendix
We introduce the following content in the appendix: SMPL-X optimization and mesh optimization details, experimental environment, network structures, limitations, and more visualizations.
SMPL-X optimization and Mesh optimization
@@
During the optimization, we render the side views and compute the side view mask loss every 10 iterations among the total 100 iterations. We show the visualization of our side view mask rendered from iteratively reconstructed Gaussisnas by HGM, initial SMPL-X, refined SMPL-X, and our finial-extracted meshes in Fig. 10, as shown in the figure, the side view masks effectively help refine the initial SMPL-X error for accurate reconstruction. For mesh refinement, we minimize the $\mathcal{L}_{1}$ loss between the predicted normal map and the rendered normal map. We also add Laplacian loss for the preservation of the local structure. We optimize 200 iterations in total. The whole Gaussian reconstruction takes around 40s and the mesh extraction takes another 30s.
Experimental environment
@@
Network structures
Our UNet model consists of 6 down blocks, 1 middle block and 5 up blocks, with an input image size of 512×512 and an output Gaussian feature map size of 256×256. We use 2 input views, resulting in a total of 256×256×4 = 131,072 output Gaussians. Each block contains several residual layers and an optional down-sample or up-sample layer. For the last 3 down blocks, the middle block, and the first 3 up blocks, we insert cross-view self-attention layers after the residual layers. The final feature maps are processed by a 1×1 convolution layer to produce 14-channel pixel-wise Gaussian features. We adopt SiLU activation and group normalization for the UNet. Our $\operatorname{Tr}_{mix}$ share the same structure, consisting of multiple self-attention blocks. Specifically, they each have 5 up blocks and 5 down blocks. The down-sample channels are [64, 128, 256, 512, 1024], and the up-sample channels are [1024, 512, 256, 128, 64]. The input dimensions for $\operatorname{Tr}_{mix}$ is 256, respectively. The output dimension for both is 14, which matches the dimension of the Gaussian features. For the attention blocks, we use a memory-efficient attention implementation.
Figure 9: Novel view synthesis comparison SHERF on HuMMan dataset.
Limitations
Currently, our method is hard to generate high-quality hands and faces, which could potentially be solved by utilizing the SMPL-X model and regional diffusion guidance such as SDS loss for further refinement.
Figure 10: SMPL-X refinment visualizaiton.
Additional results
We provide visual comparison with the SOTA NeRF-based method: SHERF [11] shown in 9. While SHERF predicts blurry results and loses fidelity, our method preserves high-frequency details and generates realistic back views such as wrinkles and hair that fit well to the front views. We also provide more mesh reconstruction results in Fig. 11.
Figure 11: Mesh reconstruction visualization for THuman 2.1 data and in the wild images.

### v5

@@ 3.4 Back-view Refinement
Back-view hallucination poses a significant challenge in single-view human reconstruction. As shown in Fig. 4, directly using diffusion models to generate a back-view image can result in incorrect perspective projection with the front-view image as well as unrealistic texture by (Ho et al., 2024). The reason is that their diffusion is conditioned on the back-view mask, and thus can only be applied for orthogonal projection where back-view mask can be directly flipped with the front-view image. However, the back-view mask is not available during our inference stage since our prediction is based on perspective projection. To address this issue, we adopt a generate-then-refine strategy that leverages the diffusion prior to produce a perspective-fitted and realistic back-view image that is suitable for the subsequent two-view fusion stage. We train a ControlNet (Zhang et al., 2023b) to give realistic details based on our coarse results as shown in the right of Fig. 3. Specifically, we generate coarse back-view rendering by our HGM for the training dataset and only train the ControlNet and keep the base Stable Diffusion model as fixed. The ControlNet loss is given as:
$\mathcal{L}_{CN}=\mathbb{E}_{\mathbf{z}_{0},t,\mathbf{c}_{t},\bm{\epsilon}_{t}\sim\mathcal{N}(0,1)}\left[\left\|\bm{\epsilon}-\epsilon_{\theta}(\mathbf{z}_{t},t,\mathbf{c}_{t},\mathbf{y})\right\|_{2}^{2}\right],$ (4)
where $y$ is the text prompt. We set it as ‘Best quality’ and the negative prompt as ‘blur, bad anatomy, bad hands, cropped, worst quality’ during inference. $\mathbf{c}_{t}$ is the coarse back-view image from our HGM rendering $I_{B}^{c}$ , which is the ControlNet condition. We carefully design the reversing process by adding small amount of noise to the VAE encoded latent of $I_{B}^{c}$ to keep the original content as much as possible. The sampling process takes around 2 seconds. In Fig. 4, we show our generated and refined back-view results, comparing them with the back-view hallucination diffusion network in SiTH (Zhang et al., 2024). Our results maintain high resolution and generate details such as the hair of the first woman and the wrinkles in the clothes. In comparison, SiTH (Zhang et al., 2024) produces artifacts and unrealistic hallucination results. Moreover, the results are also not fitted perspectively to the input image.
@@ 3.5 Two-view Reconstruction
Specifically, our fusion HGM model retains the design of the coarse HGM model with the additional refined back-view image as the input. The coarse HGM and fusion HGM models are trained separately with ground-truth one view and two views as input, as well as ground-truth SMPL-X. The objective function for HGM training includes $\mathcal{L}_{2}$ color loss, $\mathcal{L}_{rgb}$, VGG-based LPIPS perceptual loss, $\mathcal{L}_{lpips}$ (Zhang et al., 2018), and $\mathcal{L}_{2}$ background mask loss $\mathcal{L}_{bg}$ with ground truth masks. Each of these losses has corresponding weights that are treated as hyperparameters:
@@ 3.6 Implementation
Our model is trained on 4 NVIDIA RTX A6000 with batch size of 4 for 20 hours. Our input image size is 512×512 and the number of Gaussians for each view is 256×256, with a total of 65,536 Gaussians per view. For SMPL-X estimation, we use PIXIE(Feng et al., 2021). Network structures and more implementation details are in the appendix.
@@ 4 Experiments
We conduct experiments on the publicly available 3D human datasets THuman2.0 (Yu et al., 2021), CustomHumans(Ho et al., 2023) and HuMMan (Cai et al., 2022). Our method is compared with state-of-the-art (SOTA) methods in both novel view synthesis and 3D mesh reconstruction. We train our HGM on 500 human scans from THuman2.0 dataset following Zhang et al. (2024). We render the images with resolution of 512×512 and using weak perspective camera on 12 fixed cameras evenly distributed with the azimuths from 0 to 360 degree. During evaluation, all the methods are tested without the ground truth SMPL-X. We follow the train and test list from SIFU (Zhang et al., 2024) and SHERF (Hu et al., 2023) to evaluate our method on THuman2.0 and HuMMan dataset. For CustomHumans dataset we use 45 scans for cross-dataset evaluation containing loosing clothes and challenging poses. For novel view synthesis, We use PSNR, SSIM, LPIPS as evaluation metrics. For 3D reconstruction, we use commonly used Chamfer Distance (CD), Point-to-Surface Distance (P2S), and Normal Consistency as the evaluation metrics.
@@ 4.1 Novel View Synthesis
GTA Zhang et al. (2023c) 19.09 0.882 0.113 19.59 0.887 0.125
SiTH Ho et al. (2024) 17.12 0.843 0.155 18.09 0.856 0.144
LGM Tang et al. (2024a) 18.34 0.856 0.134 19.87 0.877 0.132
SV3D Voleti et al. (2024) 19.11 0.892 0.117 20.86 0.902 0.112
SIFU Zhang et al. (2024) 22.10 0.924 0.0794 20.83 0.898 0.117
@@
For novel view synthesis, we compare our method with mesh SOTA human reconstruction methods GTA Zhang et al. (2023c), ECON Xiu et al. (2023), SIFU Zhang et al. (2024) and SiTH (Ho et al., 2024), as well as multiview diffusion reconstruction method LGM(finetuned with the same training data) Tang et al. (2024a) and video diffusion method SV3D Voleti et al. (2024) on THuman2.0 Yu et al. (2021) and CustomHuman Ho et al. (2023). We also compare our method with state-of-the-art HumanNeRF methods: SHERF (Hu et al., 2023), MPS-NeRF (Gao et al., 2022), and NHP (Kwon et al., 2021) on the HuMMan dataset (Cai et al., 2022).
As shown in the Tab. 1 and Tab. 2, our method significantly surpasses state-of-the-art single-view human reconstruction methods in all evaluation metrics for the three datasets. As shown in Fig. 6, LGM Tang et al. (2024a) generates incorrect blue color and inconsistent content. SiTH Ho et al. (2024) fails to model loose clothes due to the high dependency of the SMPL-X model. Side views are blurry and unrealistic in SIFU’s results. SV3D Voleti et al. (2024) generates strange colors and wrong human pose. Compared with these methods, ours generates more realistic and consistent rendering especially for the unseen regions such as clothes wrinkles and hair that are well-fitted to the front views and more robust to initial SMPL-X estimation errors thanks to our iterative refinement. We provide rendering videos in 360 degree comparison with other methods in the Appendix. We also provide a visual comparison with the SOTA NeRF-based method SHERF Hu et al. (2023) on the HuMMan dataset in the appendix.
@@
NHP (Kwon et al., 2021) 18.99 0.845 0.182
MPS-NERF (Gao et al., 2022) 17.44 0.824 0.193
SHERF (Hu et al., 2023) 20.83 0.891 0.125
@@
For mesh reconstruction we extract the 3D mesh by densely rendering the depth map with Gaussian render and using TSDFusion to extract the surface followed by a fast optimizaiton based on the normal map obtained in section3.3.2. We compare our results with SOTA human surface reconstruction methods GTA (Zhang et al., 2023c), ECON (Xiu et al., 2023), SIFU (Zhang et al., 2024) and SiTH Ho et al. (2024). Note that our method do not use the 3D ground truth for supervision, but can also achieve best performance compared with all those fully supervised methods. Thanks to our interative SMPL-X refinement. Our method alleviates commonly occurring problems of bent legs and incorrect postures found in previous methods, as shown in Fig.5. Our HGM model reconstructs more accurate geometry with the prior learned from our dual branch Gaussian reconstruction model as well as our innovative SMPL-X refinement. We provide a qualitative 3D reconstruction comparison in Fig.7. As shown in the figure, ECON and SIFU suffer from bending legs and wrong arms problems. SiTH generates an over-smoothed surface and missing parts. Our method can reconstruct more accurate poses while preserving geometric details.
Table 3: 3D reconstruction comparison with SOTA methods. Note that only our method is trained without 3D supervision.
THuman2.0 CustomHumans
Method Chamfer $\downarrow$ P2S $\downarrow$ Normal $\uparrow$ Chamfer $\downarrow$ P2S $\downarrow$ NC $\uparrow$
ECON Xiu et al. (2023) 2.342 2.431 0.765 2.107 2.355 0.771
GTA Zhang et al. (2023c) 2.201 2.314 0.773 1.987 2.115 0.769
SiTH Ho et al. (2024) 2.519 2.442 0.786 2.223 2.584 0.785
SIFU Zhang et al. (2024) 2.063 2.205 0.792 1.864 1.976 0.778
Ours 2.134 2.118 0.823 1.729 1.835 0.834
@@ 4.3 Ablation Studies
We conduct ablation studies to evaluate the effectiveness of our SMPL-X dual branch Gaussian prediction model, coarse-to-fine refinement strategy, back-view refine ControlNet, and our SMPL-X refinement. We show the quantitative ablation results in Table 4. The performance decreases when any component is removed. SMPL-X dual branch plays an important role in adding human priors through structural features to Image Gaussians predicted by UNet. We visualize the rendered images using our model without SMPL-X dual branch as the guidance and those produced by our full model, as shown in Fig. 8. Without SMPL-X dual branch as guidance, the side-view images exhibit significant artifacts, such as misaligned arms and unnatural shapes of clothes and heads, highlighted in the red boxes. This demonstrates the effectiveness of our SMPL-X dual branch Gaussian prediction design.The predicted Gaussians lack human shape and pose priors without SMPL-X guidance, resulting in unnatural shapes and poses. Since the initial SMPL-X prediction is not accurate, we ablate the effectiveness of our iterative SMPL-X refinement on the 3D prior learned in our HGM model. We also visualize the SMPL-X refinement in the appendix. Two-view refinement doubles the number of Gaussians to improve the reconstruction quality. Gaussians tend to concentrate more on the front view without the two-view refinement strategy, leading to poorer rendering of the back part. Additionally, the diffusion-based refinement is crucial for improving the novel view synthesis quality, especially for the back-view images as shown in Fig. 4. The best performance is achieved with all three components.
@@
4.4 Discussions
Our method improves upon previous approaches like ECON Xiu et al. (2023) and SIFU Zhang et al. (2024) by leveraging our pre-trained HGM model to incorporate 3D priors, specifically side view masks, for enhanced SMPL-X refinement. The refined SMPL-X also benefits our human Gaussians reconstruction by providing structural information encoded in learnable tokens. Our 3D Gaussians-based method offers significant advantages in rendering speed, achieving 300 FPS compared to NeRF-based methods like SHERF Hu et al. (2023), which manages only 2 FPS. Furthermore, the mesh extracted from our 3D Gaussians with normal refinement attains high 3D reconstruction quality. In summary, our method excels in both high-quality rendering and accurate 3D reconstruction, offering a comprehensive solution.
@@ 5 Conclusion
Acknowledgement.
This research / project is supported by the National Research Foundation (NRF) Singapore, under its NRF-Investigatorship Programme (Award ID. NRF-NRFI09-0008).
@@ Appendix A Appendix
We introduce the following content in the appendix: SMPL-X optimization and mesh optimization details, SMPL-X evaluation, back-view details and evaluation, additional comparison, experimental environment, network structures, limitations, and more visualizations.
SMPL-X optimization and mesh optimization details.
@@
During optimization, we render SMPL-X side views and compute the side-view mask loss and normal loss for a total of 45 iterations. SMPL-X parameters are updated at each iteration. The updated SMPL-X parameters are fed into HGM to update GS every 15 iterations (3 times in total), reducing the overall optimization time. The initial SMPL-X are not input to HGM for GS prediction until after the first 15 iterations because poorly-aligned SMPL-X can lead to degraded GS. For the front view, we utilize the original image mask instead of the one rendered from GS to stabilize the process. We simultaneously render 12 views (one front-view and all the other views are considered as side views) to compute the mask loss. The loss weights are set as follows: $\mathcal{\lambda}_{front}=10$, $\mathcal{\lambda}_{side}=1$, and $\mathcal{\lambda}_{n}=0.5$. For the normal loss, we only utilize the front and back views with a pre-trained normal estimator from ICON. Throughout the optimization process, HGM remains fixed while only SMPL-X parameters are updated. The elegance of our method lies in its iterative nature: GS refines SMPL-X and better-aligned SMPL-X estimates feed back into the HGM model to generate improved 3D Gaussians, which in turn enhance the reconstruction. We show the visualization of our side-view mask rendered from iteratively reconstructed Gaussisnas by HGM, initial SMPL-X, refined SMPL-X, and our finial-extracted meshes in Fig. 11. As shown in the figure, the side view masks effectively help refine the initial SMPL-X error for accurate reconstruction. For mesh refinement, we minimize the $\mathcal{L}_{1}$ loss between the predicted normal map and the rendered normal map. We also add Laplacian loss for the preservation of the local structure.
Additional comparison with TeCH.
We compare our method with TeCH on the Customhumans dataset quantitatively in Tab. 5. TeCH needs 4-5 hours for each sample, so we use 10 samples from the CustomHumans dataset for comparison. TeCH Huang et al. (2024) has several obvious limitations compared with ours: 1) The geometry refinement from SDS is not stable and the surface is broken as shown in the left part of Fig 9 even though capturing more high-frequency geometric details. 2) Slow optimization: It needs 4-5 hours optimization, while ours use only around 90s. 3) The caption guidance can sometimes be incorrect. For example, as shown in the left part of Fig 9, the wrong caption of the gender resulted in wrong face reconstruction. 4) SMPL-X error leading to bending legs and wrong geometry, which is the same issue in SIFU, SiTH, ECON and GTA as shown in SIFU, SiTH, ECON and GTA.
Table 5: Additional evaluation with TeCH.
PSNR($\uparrow$) SSIM($\uparrow$) LPIPS($\downarrow$) CD($\downarrow$) P2S($\downarrow$) NC($\uparrow$)
Ours 24.56 0.949 0.051 1.715 1.844 0.833
TeCH 23.87 0.927 0.079 2.232 2.432 0.778
SMPL-X refinement evaluation.
We evaluate using SMPL-X initializations from PyMAF-X Zhang et al. (2023a) and PIXIE Feng et al. (2021). We compute the MPJPE (mm) using the first 22 body joints defined in SMPL-X on both THuman2.0 Yu et al. (2021) and CustomHumans Ho et al. (2023) datasets. In the Tab. 6, ‘Initial’ means the direct estimation from SMPL-X predictors. ‘w/o side-views’ represents optimization without side-views mask loss. ‘ours’ refers to our optimization with all losses, including the side-views mask loss. We can see from the table that our method successfully refines the initial SMPL-X estimates using side-view priors from our HGM, which significantly reduces the error compared with without using side-view masks. Although PyMAF-X provides better initial SMPL-X estimates than PIXIE, both methods achieve comparable MPJPE scores after optimization, as the side-view mask loss guides them toward similar convergence points. This also shows that our method is robust to diverse SMPL-X initial estimators and can effectively improve the initial SMPL-X estimation.
Table 6: SMPL-X refinement evaluation in terms of MPJPE.
PIXIE PyMAF-X
Dataset Initial($\downarrow$) w/o side-views($\downarrow$) Ours($\downarrow$) Initial($\downarrow$) w/o side-views($\downarrow$) Ours($\downarrow$)
CustomHumans 75.79 65.33 39.11 65.20 58.12 39.78
Thuman2.0 80.11 72.36 44.30 71.18 65.65 44.84
Figure 9: Visual comparison with TeCH on loose cloth and challenging pose cases.
Backview refinement details and evaluation.
We use the original ControlNet architecture and initialize the ControlNet with the ControlNet-tile model. ControlNet-tile is originally trained as an image super-resolution model. We finetune the ControlNet part with our constructed data pair at learning rate of 1e-5, with the base SD1.5 keep fixed. Data pair construction involves first training our HGM using single-view input without full convergence. Subsequently, we perform inference, downsample and render the back-view. The resulting back-view renderings, intentionally designed to have lower quality, serve as conditioning inputs for our ControlNet training. To validate the effectiveness of our proposed back-view refinement strategy, we performance quantitative evaluation with SiTH Ho et al. (2024) and  Huang et al. (2024) for back-view quality on the CustomHumans dataset. We use SSIM, LPIPS and KID as evaluation metrics between the ground truth and generated back view images. SiTH generates back-view using pure hallucination, which always generate unrealistic image as shown in Fig. 4 and Tab. 7. TeCH use SDS loss to optimize the back-view. However, the back view always fits to the wrong SMPL-X pose and imprecise text description, which leads to lower generation quality.
Table 7: Backview evaluation.
[…]
Ours 0.949 0.079 9.26
SiTH 0.855 0.123 29.8
TeCH 0.876 0.118 20.3
Figure 10: Novel view synthesis comparison SHERF on HuMMan dataset.
Experimental environment.
@@
Network structures.
Our UNet model consists of 6 down blocks, 1 middle block and 5 up blocks, with an input image size of 512×512 and an output Gaussian feature map size of 256×256. We use 2 input views, resulting in a total of $256\times 256\times 4=131,072$ output Gaussians. Each block contains several residual layers and an optional down-sample or up-sample layer. For the last 3 down blocks, the middle block, and the first 3 up blocks, we insert cross-view self-attention layers after the residual layers. The final feature maps are processed by a $1\times 1$ convolution layer to produce 14-channel pixel-wise Gaussian features. We adopt SiLU activation and group normalization for the UNet. Our $\operatorname{Tr}_{mix}$ share the same structure, consisting of multiple self-attention blocks. Specifically, they each have 5 up blocks and 5 down blocks. The down-sample channels are [64, 128, 256, 512, 1024], and the up-sample channels are [1024, 512, 256, 128, 64]. The input dimensions for $\operatorname{Tr}_{mix}$ is 256, respectively. The output dimension for both is 14, which matches the dimension of the Gaussian features. For the attention blocks, we use a memory-efficient attention implementation.
Figure 11: SMPL-X refinement visualization.
Limitations.
Currently, our method is hard to generate high-quality hands and faces, which could potentially be solved by using the SMPL-X model and regional diffusion guidance such as SDS loss for further refinement. Also, the mesh extraction process from reconstructed gaussians is not straight forward, incurring additional optimization with estimated normal map as supervision.
Additional results.
We give visual comparison with the SOTA NeRF-based method: SHERF Hu et al. (2023) shown in 10. While SHERF predicts blurry results and loses fidelity, our method preserves high-frequency details and generates realistic back views such as wrinkles and hair that fit well to the front views.